Consider a proton as a bound oscillator with a natural frequency of
4 ×10 Hz. 21 
Calculate the energy of its ground and first excited states.
1
Expert's answer
2020-04-09T09:14:24-0400

The energy levels of bound harmonic oscillator are

"(1)E_n=h\nu\cdot (n+\frac{1}{2}), n=0,1,2,..."

The ground state must correspond to a stable particle - the proton and its rest energy. Proton has rest energy "\u03f5=938MeV" This value corresponds to the oscillator frequency, which can be found from the equality of the ground state energy and the photon quantum

"\\frac12\u200bh\u03bd=\u03f5;"

"\u03bd=\\frac{2\u03f5}h\u200b=\\frac{9,38\\times10^8eV}{4.14\\times10^{\u221215}eVs}\u200b=4.53\\times10^{23}Hz"


The value of the Planck constant bar expressed in various quantities can be found in [1].

The first excited states have energy "E_n=\frac{3}{2}\cdot h\nu=3\epsilon=2.81\cdot 10^3 MeV"


Answer: The energy of ground and first excited states of proton are 938 MeV

and "2.81\times10^3MeV"